# Well-Posed Geometric Boundary data in General Relativity, III: Conformal-volume data

### v1

@@ 1. Introduction
consisting of the conformal class $[g_{{\mathcal{C}}}]$ of the induced boundary metric, together with the scalar
@@
Here $dv_{g}$ is the volume form of the ambient metric $g$ on $M$ restricted to ${\mathcal{C}}$ and $dv_{g_{{\mathcal{C}}}}$ is the volume form of the boundary metric $g_{{\mathcal{C}}}$ on ${\mathcal{C}}$. The product $\mu_{g}$ is viewed as a scalar density with respect to an arbitrarily chosen background volume form on $M$. The particular form of $\mu_{g}$ in (1.4) is chosen since it interacts well with the choice of gauge used here - the harmonic or wave coordinate gauge. Different choices of gauge may lead to other expressions for the scalar $\mu_{g}$, but this will not be pursued further here.
Note that the pair of boundary data $([g_{{\mathcal{C}}}],dv_{g_{{\mathcal{C}}}})$, when taken together, is exactly Dirichlet boundary data $g_{{\mathcal{C}}}$ studied in [3] - which was proved to be ill-posed (i.e. not well-posed) in general, i.e. without further conditions. Thus the boundary data (1.3) can be viewed as a modification of Dirichlet boundary data, involving a twisting by the volume form of $g$. The boundary data (1.3) involves no derivatives of the metric at ${\mathcal{C}}$.
@@
(1.5) $\begin{split}&\Phi:\mathbb{E}\rightarrow\mathcal{I}_{0}\times_{c}{\mathcal{B}},\\ &\Phi(g)=(g_{S},K,[g_{{\mathcal{C}}}],\mu_{g}),\\ \end{split}$
where $g_{S}$ is the metric on $S$ induced by $g$ and $K$ is the induced second fundamental form (extrinsic curvature) on $S$. The subscript $c$ denotes the compatibility conditions between the initial and boundary data at the corner $\Sigma$.
Due to the diffeomorphism invariance of the vacuum Einstein equations, it is well-known that the map $\Phi$ cannot be one-to-one and so $\Phi$ cannot be an effective parametrization of $\mathbb{E}$. Let $\mathbb{E}^{H}$ be the space of vacuum Einstein metrics in harmonic or wave gauge; cf. §2.3 for details. Let $\mathrm{Diff}_{1}(M)\subset\mathrm{Diff}_{0}(M)$ be the subgroup of diffeomorphisms $\varphi$ such that $\varphi$ is the identity to first order at $S$; equivalently,
@@
There are no $1^{\rm st}$ order restrictions for $\varphi\in\mathrm{Diff}_{0}(M)$ along ${\mathcal{C}}$. For any metric $g$ such diffeomorphisms leave the future time-like normal vector $\nu_{S}$ to $S$ in $(M,g)$ invariant: $\varphi_{*}(\nu_{S})=\nu_{S}$. It is well-known (cf. again §2.3) that for any $g\in\mathbb{E}$ there is a unique diffeomorphism $\varphi\in\mathrm{Diff}_{1}(M)$ such that
@@
The main result of this paper is the following (local-in-time) well-posedness result for the IBVP with boundary data given by (1.3). Let $\mathbb{V}_{S}^{\prime}$ be the space of smooth vector fields along $S$ nowhere tangent to $S$ (the target space for normal vector fields along $S$) and form the product space $\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$; let $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}\subset\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$ be the subset consisting of data smoothly compatible at the corner $\Sigma$, cf. §2.4 for details.
@@
The spaces $\mathbb{E}^{H}$ and $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ are smooth tame Frechet manifolds and the map
(1.7) $\begin{split}&\Phi^{H}:\mathbb{E}^{H}\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\Phi^{H}(g)=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}),\\ \end{split}$
is (locally-in-time) a smooth tame diffeomorphism of $\mathbb{E}^{H}$ onto $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ with smooth, tame inverse. Thus the IBVP with boundary data (1.3) is well-posed in harmonic gauge.
Briefly, vacuum Einstein metrics on $M$, with arbitrarily given compatible initial and boundary data, exist for a short time, and are uniquely determined in harmonic gauge by their initial data in $\mathcal{I}_{0}\times\mathbb{V}_{S}^{\prime}$ and boundary data in ${\mathcal{B}}$. Further, such solutions depend continuously and in fact smoothly on the target data in $[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$; the IBVP is stable with respect to variation of the target data. A version of Theorem 1.1 holds also off-shell (for non-vacuum Einstein metrics), cf. Theorem 6.1. We refer to [7] for details regarding Frechet spaces and tame maps.
@@
Note that this result does not depend on any choice of boundary data space ${\mathcal{B}}$; it is intrinsic to the structure of the moduli space ${\mathcal{E}}$, independent of any map to boundary data. The smoothness of the moduli space ${\mathcal{E}}$ has not been established previously. In fact it is known to be false in general for the pure Cauchy problem with compact Cauchy surface $S$ (with $\partial S=\emptyset$); this issue is closely related to the failure of the linearization stability for solutions of the Cauchy problem at Killing initial data, cf. [10].
@@
(1.8) $\begin{split}&\widehat{\Phi}:\mathbb{E}\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\widehat{\Phi}(g)=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}).\\ \end{split}$
@@
where $\pi_{*}:\mathbb{E}\rightarrow{\mathcal{E}}_{*}$ is the projection to the orbit space and ${\mathcal{E}}_{*}$ is identified with $\mathbb{E}^{H}$ as in (1.6). The map $\widehat{\Phi}$ is not constant on the fibers of the projection $\pi_{*}$; this is due to the fact that the groups $\mathrm{Diff}_{1}(M)$ and $\mathrm{Diff}_{0}(M)$ act non-trivially on the target data. To explain this in more detail, while the boundary data $[g_{{\mathcal{C}}}]$ is invariant under the action of $\mathrm{Diff}_{1}(M)$ (and $\mathrm{Diff}_{0}(M)$), the volume term $\mu_{g}$ transforms as
(1.10) $(\varphi,\mu_{g})\rightarrow\mu_{\varphi^{*}g}=(\varphi^{*}dv_{g})(dv_{g_{\mathcal{C}}})^{-2/n}.$
For example, in $4$-dimensions, the boundary degrees of freedom for the metric $g$ require $10$ independent boundary conditions, $4$ of which are associated with the choice of gauge. Here, of the remaining $6$ degrees of freedom, $5$ are geometric (corresponding to the specification of $[g_{{\mathcal{C}}}]$), but the scalar volume term $\mu_{g}$ is not (since it is not invariant under the action of $\mathrm{Diff}_{0}(M)$). This may be compared with the works in [6], where only $3$ boundary conditions are geometric (in a restricted sense); the geometric nature of the boundary conditions in [8] is less clear. Moreover, Theorem 1.1 holds in all dimensions while the works in [6] and [8] are restricted to $4$-dimensions.
A related but distinct perspective is to return to the map $\Phi^{H}$ in (1.7) and use the identification (1.6) to obtain via Theorem 1.1 the diffeomorphism
(1.11) $\begin{split}&\Phi_{*}:{\mathcal{E}}_{*}=\mathbb{E}/\mathrm{Diff}_{1}(M)\rightarrow[\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c},\\ &\Phi_{*}[g]=(g_{S},K,\nu_{S},[g_{{\mathcal{C}}}],\mu_{g}).\\ \end{split}$
@@
(1.12) $\begin{split}&\Phi_{*}:{\mathcal{E}}=\mathbb{E}/\mathrm{Diff}_{0}(M)\rightarrow\mathcal{I}_{0}\times_{c}[{\mathcal{B}}],\\ &\Phi_{*}[g]=(g_{S},K,[g_{{\mathcal{C}}}]).\\ \end{split}$
@@
$\widetilde{\mathbb{E}}=\cup_{g\in\mathbb{E}^{H}}{\mathcal{O}}_{\mathrm{Diff}_{0}^{dv_{g}}(M)}.$
@@
We point out that the uniqueness for the gauged map $\Phi^{H}$ in (1.13) cannot be improved to the ungauged map $\Phi$ in (1.5). Thus it is not true that for general $g_{1},g_{2}\in\mathbb{E}$ (not necessarily in harmonic gauge),
@@
for some $\varphi\in\mathrm{Diff}_{0}(M)$, i.e. $\Phi$ in (1.5) cannot be one-to-one modulo isometry in $\mathrm{Diff}_{0}(M)$. To see this, suppose $g^{1},g^{2}$ are a pair of vacuum Einstein metrics with the same initial data and the same Dirichlet boundary data so that $[g^{1}_{{\mathcal{C}}}]=[g^{2}_{{\mathcal{C}}}]$ and $dv_{g^{1}_{{\mathcal{C}}}}=dv_{g^{2}_{{\mathcal{C}}}}$. As noted above, $\mathrm{Diff}_{0}(M)$ acts transitively on the space of bulk volume forms $dv_{g}$ restricted to ${\mathcal{C}}$. Hence, there exists $\psi\in\mathrm{Diff}_{0}(M)$ such that for $\widetilde{g}=\psi^{*}g^{2}$, $dv_{\widetilde{g}}=\psi^{*}dv_{g^{2}}=dv_{g^{1}}$. Since $\mathrm{Diff}_{0}(M)$ leaves any volume form $dv_{g_{{\mathcal{C}}}}$ invariant, it follows that
@@
Hence the validity of (1.14) would imply $\widetilde{g}=\varphi^{*}g^{1}$ and so $g^{2}=(\varphi\circ\psi^{-1})^{*}g^{1}$. Thus (1.14) implies uniqueness of solutions to the Dirichlet boundary value problem in ${\mathcal{E}}$. However, this was proved to be false in general, (i.e. without further conditions), in [3], [4]; it follows that (1.14) is also false in general.
@@ 2.2. Localization
(2.5) $||\widetilde{g}-g_{\alpha_{0}}||_{C^{\infty}(U)}\leq\varepsilon=\varepsilon(\lambda,g),$
@@
$||f||_{H_{\widetilde{g}}^{s}(D)}^{2}=\lambda^{2s-n}||f||_{H_{g}^{s}(D)}^{2},$
$||f||_{{\mathcal{N}}_{\widetilde{g}}^{s}(M)}^{2}=\lambda^{2s-n}||f||_{{\mathcal{N}}_{g}^{s}(M)}^{2}.$
@@
(2.15) $\begin{split}&C(Q,\gamma_{S},\kappa,V_{S},\partial_{t}V_{S},\nu_{S},\gamma_{{\mathcal{C}}},V_{{\mathcal{C}}})=\begin{cases}|\kappa|^{2}-(\mathrm{tr}_{\gamma_{S}}\kappa)^{2}-R_{\gamma_{S}}-\mathrm{tr}_{g_{S}}(Q-\delta^{*}V)+3(Q(\nu_{S},\nu_{S})-\delta^{*}V(\nu_{S},\nu_{S})),\\ {\rm div_{\gamma_{S}}}\kappa-d_{S}(\mathrm{tr}_{\gamma_{S}}\kappa)-Q(\nu_{S},\cdot)+\delta^{*}V(\nu_{S},\cdot).\\ \end{cases}\end{split}$
@@
(2.24) $|v||_{\bar{\mathcal{H}}^{s}(S_{t})}^{2}\leq C[||v||_{H^{s}(S_{0})}^{2}+||\partial_{t}v||_{H^{s-1}(S_{0})}^{2}+||\nu_{{\mathcal{C}}}(v)||_{H^{s-1/3}({\mathcal{C}}_{t})}^{2}+||\varphi||_{H^{s-1}(M_{t})}^{2}],$
so there is a ’loss’ of $2/3$ of a derivative. We refer to [9, Ch.8], and in particular to [12, Thm.9] for the proof and further details.
@@
(2.26) $-\tfrac{1}{2}(\Box V+\mathrm{Ric}_{g}(V))=\beta(Q).$
@@
(2.27) $-\tfrac{1}{2}(\Box_{g}V^{\prime}_{h}+\mathrm{Ric}_{g}(V^{\prime}_{h}))+\tfrac{1}{2}\nabla_{V}V^{\prime}_{h}=\beta(F)+\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$
@@
(2.29) $-\tfrac{1}{2}[\Box V^{\prime}_{F}+\mathrm{Ric}(V^{\prime}_{F})]+\tfrac{1}{2}\nabla_{V}V^{\prime}_{F}=\beta(F),$
@@
(2.30) $-\tfrac{1}{2}[\Box W^{\prime}_{h}+\mathrm{Ric}(W^{\prime}_{h})]+\tfrac{1}{2}\nabla_{V}W^{\prime}_{h}=\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$

### v2

@@ 1. Introduction
consisting of the conformal class $g_{{\mathcal{C}}}]$ of the induced boundary metric, together with the scalar
@@
Here $dv_{g}$ is the volume form of the ambient metric $g$ on $M$ restricted to $\mathcal{C}$} and $dv_{g_{{\mathcal{C}}}}$ is the volume form of the boundary metric $g_{{\mathcal{C}}}$ on $\mathcal{C}$}. The product $\mu_{g}$ is viewed as a scalar density with respect to an arbitrarily chosen background volume form on $M$. The particular form of $\mu_{g}$ in (1.4) is chosen since it is well adapted to the choice of gauge used here - the harmonic or wave coordinate gauge. Different choices of gauge may lead to other expressions for the scalar $\mu_{g}$, but this will not be pursued further here.
Note that the pair of boundary data $([g_{{\mathcal{C}}}],dv_{g_{{\mathcal{C}}}})$, when taken together, is exactly Dirichlet boundary data $g_{{\mathcal{C}}}$ studied in [3]. This boundary data was proved to be ill-posed (i.e. not well-posed) without further conditions; this is discussed further below. The boundary data (1.3) can be viewed as a modification of Dirichlet boundary data, involving a twisting by the volume form of $g$. Observe that the boundary data (1.3) involves no derivatives of the metric at $\mathcal{C}$}.
@@
[…]
where $g_{S}$ is the metric on $S$ induced by $g$ and $K$ is the induced second fundamental form (extrinsic curvature) on $S$. The subscript $c$ denotes the $C^{\infty}$ compatibility conditions between the initial and boundary data at the corner $\Sigma$.
Due to the diffeomorphism invariance of the vacuum Einstein equations, it is well-known that the map $\Phi$ cannot be one-to-one and so $\Phi$ cannot be an effective parametrization of the space of vacuum solutions $\mathbb{E}$. Let $\mathbb{E}^{H}$ be the space of vacuum Einstein metrics in harmonic or wave gauge; cf. §2.3 for details. Let $\mathrm{Diff}_{1}(M)\subset\mathrm{Diff}_{0}(M)$ be the subgroup of diffeomorphisms $\varphi$ such that $\varphi$ is the identity to first order at $S$; equivalently, $\varphi\in\mathrm{Diff}_{0}(M)$ and
@@
There are no $1^{\rm st}$ order restrictions for $\varphi\in\mathrm{Diff}_{0}(M)$ along $\mathcal{C}$}, except for the smooth compatibility condition at the corner $\Sigma$. For any metric $g$ such diffeomorphisms leave the future time-like normal vector $\nu_{S}$ to $S$ in $(M,g)$ invariant: $\varphi_{*}(\nu_{S})=\nu_{S}$. It is well-known (cf. again §2.3) that for any $g\in\mathbb{E}$ there is a unique diffeomorphism $\varphi\in\mathrm{Diff}_{1}(M)$ such that
@@
The main result of this paper is the following (local-in-time) well-posedness result for the IBVP with boundary data given by (1.3). Let $\mathbb{V}_{S}^{\prime}$ be the space of smooth vector fields along $S$ nowhere tangent to $S$ (the target space for normal vector fields along $S$) and form the product space $\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$; let $\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}\subset\mathcal{I}_{0}\times\mathbb{V}^{\prime}\times{\mathcal{B}}$ be the subset consisting of data smoothly compatible at the corner $\Sigma$, cf. §2.4 for details.
@@
The spaces $\mathbb{E}^{H}$ and $\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ are smooth tame Frechet manifolds and the map
[…]
is (locally-in-time) a smooth tame diffeomorphism of $\mathbb{E}^{H}$ onto $\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$ with smooth, tame inverse. Thus the IBVP with boundary data (1.3) is well-posed in harmonic gauge.
Briefly, smooth vacuum Einstein metrics on $M$, with arbitrarily given smoothly compatible initial and boundary data, exist for a short time, and are uniquely determined in harmonic gauge by their initial data in $\mathcal{I}_{0}\times\mathbb{V}_{S}^{\prime}$ and boundary data in $\mathcal{B}$}. Further, such solutions depend continuously and in fact smoothly on the target data in $\mathcal{I}_{0}\times\mathbb{V}^{\prime}_{S}\times{\mathcal{B}}]_{c}$; the IBVP is stable with respect to variation of both the initial data and the boundary data in $\mathcal{B}$}. A version of Theorem 1.1 holds also off-shell (for non-vacuum Einstein metrics), cf. Theorem 6.1. We refer to [8] for details regarding Frechet spaces and tame maps.
@@
Note that Corollary 1.2 does not depend on any choice of boundary data space $\mathcal{B}$}; it is intrinsic to the structure of the moduli space $\mathcal{E}$}, independent of any map to boundary data. The smoothness of the moduli space $\mathcal{E}$} has not been established previously. In fact it is known to be false in general for the pure Cauchy problem with compact Cauchy surface $S$ with $\partial S=\emptyset$; this issue is closely related to the failure of the linearization stability for solutions of the Cauchy problem at Killing initial data, cf. [12].
@@
[…]
@@
where $\pi_{*}:\mathbb{E}\rightarrow{\mathcal{E}}_{*}$ is the projection to the orbit space and $\mathcal{E}$$_{*}$ is identified with $\mathbb{E}^{H}$ as in (1.6). The map $\widehat{\Phi}$ is not constant on the fibers of the projection $\pi_{*}$; this is due to the fact that the groups $\mathrm{Diff}_{1}(M)$ and $\mathrm{Diff}_{0}(M)$ act non-trivially on the target data. To explain this in more detail, while the boundary data $g_{{\mathcal{C}}}]$ is invariant under the action of $\mathrm{Diff}_{1}(M)$ (and $\mathrm{Diff}_{0}(M)$), the volume term $\mu_{g}$ transforms as
(1.10) $(\varphi,\mu_{g})\rightarrow\mu_{\varphi^{*}g}=(\varphi^{*}dv_{g})(dv_{g_{\mathcal{C}}})^{-2/n},$
and generally $\varphi^{*}dv_{g}\neq dv_{g}$ along $\mathcal{C}$}.
For example, in $4$-dimensions, the boundary degrees of freedom for the metric $g$ require $10$ independent boundary conditions, $4$ of which are associated with the choice of gauge. Here, of the remaining $6$ degrees of freedom, $5$ are geometric (corresponding to the specification of $g_{{\mathcal{C}}}]$), but the scalar volume term $\mu_{g}$ is not (since it is not invariant under the action of $\mathrm{Diff}_{0}(M)$). This may be compared with the works in [7],[9]. In [7] at most $3$ boundary conditions may be considered geometric in a restricted sense while the geometric nature of the boundary conditions in [9] is less clear. Moreover, Theorem 1.1 holds in all dimensions while the works in [7] and [9] are restricted to $4$-dimensions.
Theorem 1.1 should be compared with two other recent related results. First, in [3], the IBVP was proved to be well-posed for Dirichlet boundary data in an open neighborhood of solutions $(M,g)$ for which the Brown-York stress-energy tensor
(1.11) $\Pi=Hg_{{\mathcal{C}}}-A,$
has the same Lorentz signature $(-+\cdots+)$ as the induced metric $g_{{\mathcal{C}}}$ on $\mathcal{C}$}. Here $A$ is the second fundamental form of ${\mathcal{C}}\subset(M,g)$ and $H=\mathrm{tr}_{g_{{\mathcal{C}}}}A$ is the mean curvature of the boundary. While this is the first result on well-posedness for geometric boundary data, it does not hold in full generality, i.e. without some assumption such as (1.11), cf. [4], [5] for examples of ill-posedness and further discussion. Next, it was recently proved in [10] that the IBVP for (the geometric) conformal mean curvature boundary data $([g_{{\mathcal{C}}}],H)$ is ill-posed at the linearized level at certain backgrounds. More precisely, this was proved to be the case when the IBVP is linearized at a standard spherical cylinder in Minkowski $\mathbb{R}^{1,3}$, These remarks illustrate some the difficulties in establishing a general well-posedness result for a geometric IBVP for the Einstein equations, Theorem 1.1 may be viewed as a further step in this direction.
Returning again to Theorem 1.1, consider the map $\Phi^{H}$ in (1.7) and use the identification (1.6) to obtain via Theorem 1.1 the diffeomorphism
[…]
@@
[…]
Corollary 1.3.
The map $\Phi_{*}$ in (1.13) is surjective. Thus, given arbitrary smooth initial data $(\gamma_{S},\kappa)\in\mathcal{I}_{0}$ together with a smoothly compatible conformal class $\gamma_{{\mathcal{C}}}]\in{\rm Conf}({\mathcal{C}})$, there exist smooth Einstein metrics $(M,g)$ such that $\Phi_{*}(g)=(g_{S},K,[g_{{\mathcal{C}}}])=(\gamma_{S},\kappa,[\gamma_{{\mathcal{C}}}])$.
However the map $\Phi_{*}$ is of course no longer injective. There is a full scalar degree of freedom in the domain $\mathcal{E}$} mapping to the same point in target space. Theorem 1.1 shows that the fiber $(\Phi_{*})^{-1}(\gamma_{S},\kappa,[\gamma_{{\mathcal{C}}}])$ in $\mathcal{E}$} is naturally and smoothly identified with $\{\mu\}=C^{\infty}({\mathcal{C}})$. This is less clear when $\mu$ is replaced by the mean curvature $H$ of the boundary.
@@
$\widetilde{\mathbb{E}}=\bigcup_{g\in\mathbb{E}^{H}}{\mathcal{O}}_{\mathrm{Diff}_{0}^{dv_{g}}(M)}.$
@@
We point out that this uniqueness for the gauged map $\Phi^{H}$ in (1.14) cannot be improved to the ungauged map $\Phi$ in (1.5). Thus it is not true that for general $g_{1},g_{2}\in\mathbb{E}$ (not necessarily in harmonic gauge),
@@
for some $\varphi\in\mathrm{Diff}_{0}(M)$, i.e. $\Phi$ in (1.5) cannot be one-to-one modulo isometry in $\mathrm{Diff}_{0}(M)$. To see this, suppose $g^{1},g^{2}$ are a pair of vacuum Einstein metrics with the same initial data and the same Dirichlet boundary data so that $g^{1}_{{\mathcal{C}}}]=[g^{2}_{{\mathcal{C}}}]$ and $dv_{g^{1}_{{\mathcal{C}}}}=dv_{g^{2}_{{\mathcal{C}}}}$. As noted above, $\mathrm{Diff}_{0}(M)$ acts transitively on the space of bulk volume forms $dv_{g}$ restricted to $\mathcal{C}$}. Hence, there exists $\psi\in\mathrm{Diff}_{0}(M)$ such that for $\widetilde{g}=\psi^{*}g^{2}$, $dv_{\widetilde{g}}=\psi^{*}dv_{g^{2}}=dv_{g^{1}}$. Since $\mathrm{Diff}_{0}(M)$ leaves any volume form $dv_{g_{{\mathcal{C}}}}$ invariant, it follows that
@@
Hence the validity of (1.15) would imply $\widetilde{g}=\varphi^{*}g^{1}$ and so $g^{2}=(\varphi\circ\psi^{-1})^{*}g^{1}$. Thus (1.15) implies uniqueness of solutions to the Dirichlet boundary value problem in $\mathcal{E}$}. However, this was proved to be false in general, (i.e. without further conditions), in [3], [4], [5]; it follows that (1.15) is also false in general.
@@ 2.2. Localization
(2.5) $\widetilde{g}-g_{\alpha_{0}}||_{C^{\infty}(U)}\leq\varepsilon=\varepsilon(\lambda,g),$
@@
[…]
@@
[…]
@@
[…]
so there is a ’loss’ of $2/3$ of a derivative. We refer to [11, Ch.8], and in particular to [14, Thm.9] for the proof and further details.
@@
(2.26) $\tfrac{1}{2}(\Box V+\mathrm{Ric}_{g}(V))=\beta(Q).$
@@
(2.27) $\tfrac{1}{2}(\Box_{g}V^{\prime}_{h}+\mathrm{Ric}_{g}(V^{\prime}_{h}))+\tfrac{1}{2}\nabla_{V}V^{\prime}_{h}=\beta(F)+\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$
@@
(2.29) $\tfrac{1}{2}[\Box V^{\prime}_{F}+\mathrm{Ric}(V^{\prime}_{F})]+\tfrac{1}{2}\nabla_{V}V^{\prime}_{F}=\beta(F),$
@@
(2.30) $\tfrac{1}{2}[\Box W^{\prime}_{h}+\mathrm{Ric}(W^{\prime}_{h})]+\tfrac{1}{2}\nabla_{V}W^{\prime}_{h}=\beta^{\prime}_{h}\mathrm{Ric}+O_{2,1}(V,h),$